# DreamO: A Unified Framework for Image Customization

### v3

@@ 3.3 Routing Constraint
$L=\lambda_{diff}\cdot L_{diff}+\lambda_{route}\cdot L_{route}+\lambda_{holder}\cdot L_{holder},$ (5)
@@
where $\lambda_{diff}$, $\lambda_{route}$ and $\lambda_{holder}$ are loss weights. Note that in order to allow the model to accommodate regular text input, we introduce normal text without placeholders with a probability of $50\%$ during training and discard $L_{holder}$ accordingly.
As shown in the second row of Fig. 2, after training with routing constraint, the attention of the condition image clearly focuses on the target subject, and the generated result shows improved consistency with the reference image in terms of details. In addition to the improved consistency, this strategy also helps the decoupling in multi-reference cases. More details are presented in the ablation study, i.e., Fig. 10. Fig. 3 shows the effect of placeholder, which can control the placement of conditions in the generated result.
3.4 Training Data Construction
@@
In this work, we aim to achieve generalized image customization. To this end, we collect a training dataset covering a wide range of tasks.
3.4.1 Identity Paired Data
@@
Since high-quality identity paired data [26] is difficult to collect from the Internet, we adopt the open-source ID customization method PuLID [10] for dataset construction, as we find it exhibits high facial similarity and strong prompt alignment. Specifically, we provide PuLID-SDXL with a reference face image and a text prompt describing the desired style. This allows PuLID-SDXL to control the style of the generated portrait, resulting in training pairs of the form (reference face, prompt, stylized face). For photo-realistic scenes, we generate two images of the same identity using PuLID-FLUX, which then serve as mutual references.
3.4.2 Subject-driven Data
@@
To enhance the subject-driven generation of human references, we build another dataset with a pipeline like MoiveGen [37]. Starting from a dataset of long videos, content-aware scene detection is used to get multiple short video clips. To get the inner-clip instance matching, we use Mask2Former [5] to get the mask of all the people in the key frames and do object tracking in the video clip. To get the cross-clip instance matching, we adopt SigLip [58] embedding of all human objects and do clustering. After that, we use GPT-4o [19] with image and colorful instances-mask input to get the instance grounded caption.
Figure 4: Visualization of cross-attention maps in some unseen scenarios.
3.4.3 Try-on Data
In this task, we construct a paired try-on dataset using two sources. A portion of the data is collected directly from the web as paired images of models and clothing. For the other part, we first crawl high-quality model images as ground truth, then apply image segmentation to extract the clothing items and construct the corresponding paired data. All collected images undergo manual filtering to remove low-quality samples. Clothing regions are extracted using a segmentation algorithm [20, 21].
3.4.4 Style-driven Data
In this paper, we aim to address two types of style transfer tasks: (1) style reference image control with text description of content, and (2) style reference image with content reference image. For the first type of style transfer task, the training data needs to include reference images of the same style along with target images. As shown in Fig. 12, we employ an internal style transfer model (i.e., an SDXL-based model utilizing a decoupling strategy similar to InstantStyle [49] to generate images of the same style but different content under the guidance of two distinct prompts. For the second type of task, the training requires style reference images, content reference images, and target images. Here, the target images share the same style as the style reference, while maintaining the same content structure as the content reference image. Based on the Type 1 training data, we construct content reference images by generating natural images corresponding to style images through Canny-guided Flux [1]. More details are presented in A.1.
3.4.5 Routing Mask Extraction
To obtain the labels for the routing constraint (i.e., Eq. 3), we extract target subject masks from various types of training data. Specifically, for the X2I-Subject [54] dataset, we use the InternVL [4] model to extract subject descriptions, then employ LISA [24] with the descriptions and target images to generate subject masks. For the Subject200K [47] dataset, InternVL is used to obtain subject names, followed by LISA to predict the corresponding masks. For face data, LISA is directly used to extract face masks from target images. More details are presented in A.1.
Although the routing constraint is applied on a limited set of training data (i.e., Subject200K [47], face data, and X2I-Subject [54]), our method demonstrates strong generalization to unseen scenarios. As illustrated in Fig. 4, the first row shows that in the try-on task, the condition image accurately localizes the target region. In the second row, under more complex conditions, the model effectively distinguishes the regions associated with different condition images. Specifically, the identity reference image focuses on the face and body, while the image with sunglasses attends directly to the sunglasses region, avoiding the face.
Figure 5: The progressive training pipeline of our method. Left column shows the three training stages of our method. Right column shows the generation capability after the training of each stage.
3.5 Progressive Training Process
During the experiment, we find that training directly on all the data makes convergence difficult. This is primarily due to the limited capacity of the optimization parameters, which makes it challenging for the model to capture task-specific capabilities under complex data distributions. Furthermore, due to the impact of the image quality of the training data, there is a divergence from the generative prior of Flux [1] in the quality of the generation after training.
To address these issues, we design a progressive training strategy that allows the model to smoothly converge across different tasks while rectifying the impact of training data on the high-quality generation prior of Flux [1]. The training pipeline is presented in Fig. 5. Specifically, we first optimize the model on subject-driven training data to initiate the model with a consistency-preserving capability. Note that the Subject200K [47] dataset used in the training is generated by the base model (i.e., Flux), and thus shares a similar distribution with the model generation space, which facilitates fast convergence. Since the X2I-subject [54] dataset is synthetically generated by MS-Diffusion [51], a lot of training samples contain undesired artifacts and distortions. Therefore, during this warm-up stage, the two-column Subject200K images described in Sec. 3.4.2 are also utilized as part of the training data to facilitate rapid convergence of the multi-subject generation control. The right part of Fig. 5 illustrates that after the first training stage, the model acquires an initial subject-driven generation capability and presents strong text-following performance. In the second training stage, we incorporate all the training data and perform full-data tuning. This allows the model to further converge on all subtasks defined in this work.
4 Experiment
4.1 Implementation Details
@@
In this paper, we adopt Flux-1.0-dev as our base model. The rank of additional LoRA [43] is set as 128, resulting in a total parameter increase of 707M. For the training process, we employ the Adam [28] optimizer with a learning rate of 4e-5 and train on 8 NVIDIA A100 80G GPUs. The batch size is set as 8. The first training stage consists of 20K iterations, followed by 90K iterations in the second stage, and finally 3K iterations in the last training stage.
4.2 Model Capabilities
Our proposed DreamO is a unified framework capable of handling a wide range of image customization tasks. Fig. 6 illustrates identity-driven image generation, where our method allows for both individual-specific customization and compositional identity control. That is, the model can generate images that preserve the identity of a single person or blend features from multiple individuals, while allowing precise control over other attributes and scene details through text input. Fig. 7 shows the results of object-driven image customization. One can see that, the proposed DreamO supports both single-subject conditioning and multi-subject composition, demonstrating the ability to incorporate diverse visual elements into a unified output while maintaining semantic consistency. Fig. 8 demonstrates our try-on capability. This includes visual try-on for specific identities and imaginative applications to arbitrary subjects, enabling creative reinterpretation of garments and their features. Fig. 9 demonstrates style-driven customization, with style exclusively guided by reference images, and other attributes controlled via text or image inputs, enabling fine-grained multimodal manipulation. The above results further highlight our method’s ability to integrate heterogeneous control signals—such as identity, object, and try-on, enabling more expressive and complex image customization.
Figure 6: The capability of our proposed DreamO in identity-driven image customization.
Figure 7: The capability of our proposed DreamO in subject-driven image customization.
Figure 8: The capability of our proposed DreamO in try-on image customization.
Figure 9: The capability of our proposed DreamO in style-driven image customization.
4.3 Ablation Study
@@
Progressive training. To enable the model to better converge on all sub-tasks under complex data distributions and to rectify the impact of training data distribution on generation quality, we design a progressive training strategy. The effectiveness of this strategy is demonstrated in Fig. 11. One can see that directly training the model on all datasets leads to suboptimal convergence, particularly in complex tasks such as multi-subject consistency. Warming up on a smaller and easier-to-learn dataset (e.g., Subject200K [47]) before joint training improves convergence, but the generation quality is easily influenced by the training data distribution, deviating from the generation priors of Flux. By introducing an image quality tuning stage, the model can produce higher-quality generation results.
5 Conclusion
@@
In this work, we present DreamO, a unified framework for generalized image customization across various condition types (e.g., identity, style, subject, and try-on) within a single pre-trained DiT framework. We first construct a large-scale training dataset encompassing a wide range of image customization tasks. By integrating all condition types into the DiT input sequence and introducing a feature routing constraint, DreamO ensures high-fidelity consistency while effectively disentangling heterogeneous control signals. To address the challenges of learning strong task-specific capabilities under complex data distributions while preserving the generative priors of the base model (i.e., Flux), we design a progressive training strategy. This approach enables the model to gradually acquire diverse control capabilities while maintaining the image quality of the base model. Extensive experiments demonstrate that DreamO can perform various image customization tasks with high quality. Furthermore, its lightweight LoRA-based design allows for efficient deployment with low computational cost.
Figure 10: The ablation study of routing constraint in our proposed DreamO.
Figure 11: The ablation study of progressive training in our proposed DreamO.
Appendix A Appendix
A.1 Data Construction
Style-driven data. The pipeline for constructing style transfer training data in this paper is shown in Fig. 12. First, we employ an internal style transfer model to generate paired images with the same style but different content, using different prompts under the same style reference. These paired images form the style transfer training data driven by the style reference image. Then, we use Canny-guided Flux [1] to generate a natural image with the same content by referencing the edge of one of the style images, thereby forming a style-reference and content-reference-driven style transfer training dataset with two style images.
Figure 12: The pipeline for data construction of style transfer. (Style 1, Style 2) constitutes a stylization training pair guided by style reference images. (Style 1, Style 2, Content Image) constitutes a stylization training triplet guided by both style reference images and a content reference image.
Routing mask extraction. The pipeline for constructing style transfer training data in this paper is shown in Fig. 13. First, for the multi-subject dataset X2I-Subject [54], we utilize the VLM model InternVL [4] to extract descriptions of each target subject. Then, we use LISA [24] with the subject descriptions and target images as input to extract the masks of the target subjects in the target images. For the Subject200K dataset [47], we employ InternVL to obtain the name of the subject, followed by LISA to predict the mask of the subject with the name in the target image. For face data, we directly use LISA to extract face masks from target images.
Figure 13: The pipeline for subject mask extraction from target images. The processing data includes X2I-Subject [54], Subject200K [47], and face data.

### v4

@@ 3.3. Routing Constraint
(5) $L=\lambda_{diff}\cdot L_{diff}+\lambda_{route}\cdot L_{route}+\lambda_{holder}\cdot L_{holder},$
@@
where $\lambda_{diff}$, $\lambda_{route}$ and $\lambda_{holder}$ are loss weights. To allow the model to handle regular text input, we introduce normal text without placeholders with a probability of $50\%$ and discard $L_{holder}$ accordingly. Note that $L_{route}$ and $L_{holder}$ do not incur significant additional computational overhead during training (2.5s/iter vs. 3s/iter).
Figure 4. The progressive training pipeline of our method. Left column shows the three training stages of our method. Right column shows the generation capability after the training of each stage.
Figure 5. Visual comparison between our DreamO and other methods.
3.4. Training Data Construction
@@
To achieve generalized image customization, we collect training data, covering a wide range of tasks.
@@
Identity paired data. Since high-quality identity paired data (Li et al., 2024) is difficult to collect from the Internet, we adopt the open-source ID customization method PuLID (Guo et al., 2024) for dataset construction. Specifically, we generate two images of the same identity using PuLID-FLUX, which then serve as mutual references. We also provide PuLID-SDXL with a reference face image and a text prompt describing the desired style to produce stylized training pairs. Finally, we collect 150K photorealistic data and 60K stylized identity data.
Subject-driven data. For single-subject-driven image customization, we utilize the Subject200K (Tan et al., 2024) dataset as part of training data. To rectify the absence of character-related conditions, we collect 100K paired character-related data through retrieval. For multi-subject-driven image customization, we construct some two-column images through concatenation on the Subject200K dataset. In addition, we employ X2I-subject (Xiao et al., 2024a) dataset in multi-subject-driven training. To improve human-driven generation, we develop a pipeline similar to MovieGen (Polyak et al., 2025). Starting with a long-video dataset, we apply content-aware scene detection to extract short clips. Mask2Former (Cheng et al., 2022) is used to generate human masks for key frames and perform object tracking. For cross-clip instance matching, we use SigLip (Zhai et al., 2023) embeddings and clustering.
Try-on data. We create a paired try-on dataset from two sources. One part consists of paired model and clothing images collected directly from the Web. For the other, we first crawl high-quality model images as ground truth, then use image segmentation (Jin, 2023; Jin et al., 2024) to extract clothing items and generate the corresponding pairs. All images are manually filtered to remove low-quality samples, resulting in a dataset of 500K try-on pairs.
Style-driven data. This paper tackles two style transfer tasks: (1) style reference control with text descriptions of content, and (2) style reference with content reference images. For the first task, we use an internal style customization model based on SDXL to generate images with the same style but different content from two distinct prompts. For the second task, training requires style reference images, content reference images, and target images, where the target image must match both the style and content structure of the reference images. Based on the type 1 dataset, we produce the content reference for each style image using Canny-guided Flux (flu, 2023c). More details are provided in the Appendix.
Routing mask extraction. To obtain the labels for the routing constraint (i.e., Eq.3), we use LISA(Lai et al., 2024) to extract mask of object conditioned on the text descriptions. In certain complex datasets, we employ InternVL (Chen et al., 2024a) to generate description of the target object. More details are provided in the Appendix.
Although the training data is constructed separately for different tasks, we observe the emergence of some cross-task capabilities. For instance, the model can customize the combination of ID and Try-on (as shown in Fig. 12), which does not exist in the training data.
3.5. Progressive Training Process
In experiments, we find that training directly on all data makes convergence difficult. This is mainly due to the limited capacity of LoRA (Ryu, 2023) optimization, making it difficult for the model to capture task-specific capabilities under complex data distributions. In addition to convergence, the quality of the output after training deviates from the original prior of Flux (flu, 2023a). This divergence is caused by the impact of some low-quality training samples.
Table 1. Quantitative evaluation of subject-driven customization.
Single-subject Customization Multi-subject Customization
MS-Diffusion OmniGen OminiControl DreamO MS-Diffusion OmniGen DreamO
CLIP-sim $\uparrow$ 0.8989 0.8824 0.8220 0.9150 0.7686 0.7605 0.7775
DINO-sim $\uparrow$ 0.7746 0.7582 0.6089 0.8056 0.6113 0.5646 0.6253
Text-sim $\uparrow$ 31.78 31.74 31.12 31.92 31.34 29.55 31.46
To address these issues, we design a progressive training strategy that allows the model to smoothly converge across different tasks while mitigating the influence of training data on the generation prior of Flux. The training pipeline is shown in Fig. 4. Specifically, we first optimize the model on subject-driven training data to initiate the model with a consistency-preserving capability. Note that the Subject200K (Tan et al., 2024) training data is generated by the base model (i.e., Flux), thus shares a similar distribution with the model generation space, which facilitates fast convergence. Since the X2I-subject (Xiao et al., 2024a) dataset is synthetically generated by MS-Diffusion (Wang et al., 2024b), a lot of training samples contain undesired artifacts and distortions. Therefore, during this warm-up stage, the two-column Subject200K images described in Sec. 3.4 are also utilized as part of the training data to facilitate rapid convergence of the multi-reference control. The right part of Fig. 4 illustrates that after the first training stage, the model acquires an initial subject-driven generation capability and presents strong text-following performance. In the second training stage, we incorporate all the training data and perform full-data tuning. This allows the model to further converge on all subtasks defined in this work.
@@ 3.5. Progressive Training Process
StyleAlign StyleShot InstantStyle DEADiff CSGO DreamO
Style-sim $\uparrow$ 0.7122 0.6922 0.6988 0.7269 0.7296 0.7340
Text-sim $\uparrow$ 0.2566 0.2693 0.2721 0.2656 0.2701 0.2750
4. Experiment
4.1. Implementation Details
@@
In this paper, we adopt Flux-1.0-dev as the base model. The rank of LoRA (Ryu, 2023) is set as 128, resulting in a total parameter increase of 478M. During training, we employ the Adam (Loshchilov and Hutter, 2017) optimizer with a learning rate of 4e-5 and train on 8 NVIDIA A100 80G GPUs. The batch size is set as 8. The first training stage consists of 20K iterations, followed by 90K iterations in the second stage, and finally 3K iterations in the last training stage. In inference, we use Flux-Turbo (flu, 2023b) for acceleration, enabling the generation of $1024\times 1024$ results within 10s. Unless specified, all results in this paper are based on the Turbo model.
4.2. Qualitative Comparsion
To validate the performance of DreamO, we conduct comparisons with recent state-of-the-art methods across multiple subtasks. The visual comparison is shown in Fig. 5. The first part presents a comparison between DreamO and SOTA identity-customization methods i.e., PhotoMaker (Li et al., 2024), InstantID (Wang et al., 2024a), PuLID (Guo et al., 2024). The results demonstrate that DreamO can inject identity information with high fidelity across various scenes, while offering impressive flexibility for customization.
The second part compares DreamO with recent subject-customization methods, including single-task frameworks (i.e., MS-Diffusion (Wang et al., 2024b)) and unified generation frameworks (i.e., OmniGen (Xiao et al., 2024a), OminiControl (Tan et al., 2024)). The results demonstrate that DreamO achieves higher subject fidelity and better text consistency in both single-subject and multi-subbject scenarios.
The third part shows comparison in virtual try-on, indicating that DreamO can effectively place clothing in scenes that align with the provided text with high fidelity to the reference images. Unlike methods like IMAGDressing, which produce high-fidelity clothing but lose text alignment, DreamO maintains both.
The last part presents the comparison between DreamO and recent style-customization methods, i.e., StyleShot (Gao et al., 2024), StyleAlign (Wu et al., 2021), InstantStyle (Wang et al., 2024c), DeaDiff (Qi et al., 2024), and CSGO (Xing et al., 2024). It can be observed that DreamO has weaker content intrusion, better text alignment, and higher style fidelity in the generated results.
Table 3. Quantitative evaluation of identity-driven customization.
PhotoMaker InstantID PuLID DreamO
Face-sim $\uparrow$ 0.212 0.590 0.5829 0.607
Text-sim $\uparrow$ 0.2520 0.2294 0.2534 0.2570
Table 4. Quantitative evaluation of try-on.
MagCloth IMAGDressing OmniGen OminiControl DreamO
CLIP-sim $\uparrow$ 0.5977 0.8405 0.7265 0.7065 0.7613
Text-sim $\uparrow$ 30.17 17.74 27.83 28.79 30.47
Figure 6. The ablation study of the placeholder-to-image routing constraint.
4.3. Quantitative Comparison
In addition to qualitative comparison, we conduct quantitative comparisons for each task. We present the comparison of identity customization in Tab. 3, which is evaluated on Unsplash-50 (Gal et al., 2024). Here, we provide 9 prompts for each face. Following PuLID (Guo et al., 2024), the Face-Sim represents the ID cosine similarity, with ID embeddings extracted by CurricularFace (Huang et al., 2020). We also compute the CLIP cosine similarity between the generated result and the prompt to measure the text-following ability of different methods. As can be seen, our DreamO shows better face similarity and text-based customization ability.
Tab. 1 presents the quantitative comparison in single- and multi-subject customization. We use DreamBench (Ruiz et al., 2023) as the testset of single-subject customization. For multi-subject customization, we randomly select 20 pairs from DreamBench and provide 25 prompts for each. During testing, we generate four images with different seeds for each test sample. Here, we calculate the CLIP cosine similarity and Dino (Caron et al., 2021) cosine similarity between the generated result and the reference images as a measure of subject consistency. To improve accuracy, we remove the background of the generated results and then calculate similarity. Additionally, we employ CLIP cosine similarity between the text description and the generated result as a measure of content alignment. Tab. 1 shows that our method outperforms others in subject consistency while demonstrating strong text-following ability.
The quantitative comparison of the try-on is presented in Tab. 4. We select 300 reference garments from VITON-HD (Choi et al., 2021) encompassing various styles and colors as the test data. During testing, we provide 10 prompts for each cloth. The CLIP consine similarity between the generated result and reference cloth is employed to measure the try-on accuracy. Here, we only crop the cloth from the result to compute the CLIP-sim. The CLIP cosine similarity between the generated results and the prompt is employed to measure the text-following ability of different methods. The result in Tab. 4 shows the attractive performance of our DreamO. Although IMAGDressing (Shen et al., 2025) has higher CLIP-sim, it can only generate images with a white background (i.e., Fig. 5 ) with little text-following ability.
For style customization, we construct an evaluation dataset, containing $249$ style images and $24$ prompt. Each method generates $249\times 24$ style customization results. We use the pre-trained CSD (Somepalli et al., 2024) to extract the style feature of generated results and reference images, and calculate the cosine similarity between them as a measure of style consistency. Furthermore, we compute the CLIP (Radford et al., 2021) similarity between stylized results and text descriptions as the metric of content consistency. The results are shown in Tab. 2, which demonstrate that our method has better performance in style consistency and content consistency.
Figure 7. The ablation study of routing constraint in our proposed DreamO.
Figure 8. The ablation study of progressive training in our proposed DreamO.
4.4. User Study
In addition to the automatic evaluation metrics, we also conduct a user study for manual evaluation of different methods. Specifically, for each task (i.e., style, object, identity, and try-on customization), we assigned 6 test samples and invited 20 volunteers to rate on three aspects: text alignment, reference alignment, and image quality. The scoring range was set from 0 to 5, where a higher score indicates greater satisfaction. Fig. 9 shows that our DreamO achieves better performance in these three evaluation aspects.
4.5. Ablation Study
Table 5. Ablation study of different model settings in multi-subject-driven customization.
w/o CE w/o RC w/o PT DreamO
CLIP-sim $\uparrow$ 0.7697 0.7448 0.7349 0.7775
DINO-sim $\uparrow$ 0.6097 0.5540 0.5381 0.6253
Text-sim $\uparrow$ 31.26 28.42 28.31 31.46
@@ 4.5. Ablation Study
Progressive training. To enable the model to better converge on all sub-tasks under complex data distributions and to rectify the impact of training data distribution on generation quality, we design a progressive training strategy. The effectiveness of this strategy is demonstrated in Fig. 8. One can see that directly training the model on all datasets leads to suboptimal convergence, particularly in complex tasks such as multi-subject consistency. Warming up on a smaller and easier-to-learn dataset (e.g., Subject200K (Tan et al., 2024)) before joint training improves convergence, but the generation quality is easily influenced by the training data distribution, deviating from the generation priors of Flux. By introducing an image quality tuning stage, the model can produce higher-quality generation results.
Placeholder-to-image routing constraint. As demonstrated in Eq. 4, this paper designs a placeholder-to-image routing constraint to build the routing relationship between placeholders and specific images. Fig. 6 shows the effect of this loss term. It can be seen that without this loss, placeholders struggle to precisely control their corresponding images. After applying this loss, placeholders can bind to specific reference objects, enabling individual control over particular objects during multi-subject customization.
Quantitative results. In addition to the visual comparison, we show the quantitative results of the ablation study, as shown in Tab. 5. The experiment is conducted on multi-subject-driven customization. It can be seen that not using the routing constraint (RC) and progressive training strategy (PT) significantly impacts performance, leading to a decrease in the reference consistency and text following. We also study the role of condition embedding (CE), and its absence results in a decline in the reference consistency.
5. Conclusion
@@
In this study, we introduce DreamO, a unified framework designed for generalized image customization across diverse condition types (e.g., identity, style, subject, and try-on) within a single pre-trained DiT architecture. To facilitate this, we construct a large-scale training dataset. By embedding all condition types into the DiT input sequence and incorporating a feature routing constraint, DreamO achieves high-fidelity consistency while effectively disentangling heterogeneous control signals. In addition, we design a progressive training strategy that enables the model to incrementally acquire diverse control capabilities under complex data distributions, while maintaining the image quality inherent to the base model. Comprehensive experiments demonstrate that DreamO excels in performing a wide range of image customization tasks with high-quality results.